Marc, a single taxpayer, earns $30,000 in taxable income and $1,000 in interest from an investment in city of Birmingham bonds. Using the U.S. tax rate schedule for year 2019, what is his current marginal tax rate? (Use Tax rate schedule)
Answer #1

hIs Tax Rate would br as under

on 9700 = (9700-0)X10% = 970

On Balance Rs 20300 = (30,000-9700)*12% = 2436

Total Tax = 3406

therefore, Average Tax Rate = 3406/31000 = 10.99%

And Current MArginal Tax Rate = 12%